# OV-MER: Towards Open-Vocabulary Multimodal Emotion Recognition

### v2

@@ Appendix D Related Work
MER has rapidly developed in recent years [35]. Current research mainly focuses on building more efficient architectures to achieve higher accuracy on benchmark datasets [36]. For example, [11] proposed a tensor fusion network that addressed the MER task by leveraging interactions among unimodal, bimodal, and trimodal inputs. [26] introduced a Transformer-based model that learned implicit alignment between different modalities and achieved promising results. [21] further established MERBench, involving various features, fusion strategies, and datasets. In emotion recognition, benchmark datasets usually limit the label space to basic emotions and use majority voting to determine the most likely one or more labels [9, 7]. However, emotional categories extend far beyond basic emotions. Restricting the label space will inevitably overlook some nuanced emotions. To address this issue, we extend traditional MER to OV-MER, which allows for the prediction of any number and categories of emotions.
@@
Its main goal is to identify categories beyond the annotated label space [37], which has been applied in various fields, such as object detection [38], segmentation [39], and scene understanding [40]. For example, the object detection dataset COCO [41] contains 80 categories, while objects in the real world are nearly infinite, highlighting the importance of open vocabulary learning. This paper makes the first attempt to address MER in an open-vocabulary manner. Compared to other tasks, MER is more difficult as it requires considering multimodal clues and subtle temporal variations.
Appendix E More Examples
Figures 8$\sim$10 provide more examples to visualize the difference between one-hot and OV labels. This paper uses emotion-related descriptions as a bridge for OV label extraction. Since the original video contains real people, we use DemoAI to remove personal information to address copyright concerns. Our OV-MERD dataset is derived from the MER2023 dataset [9] with further annotations. Therefore, for the original data, please download the MER2023 dataset.
@@ Appendix E More Examples
Appendix F Summary of Abbreviations
@@
Appendix G Dataset Comparison
@@
MSP-Podcast [42] A Dimensional 3 1
SST [43] T Dimensional 1 1
Cornell [44] T Dimensional 1 1
Large Movie [45] T Dimensional 1 1
ICT-MMMO [46] A,V,T Dimensional 1 1
YouTube [47] A,V,T Dimensional 1 1
MOUD [10] A,V,T Dimensional 1 1
CMU-MOSI [11] A,V,T Dimensional 1 1
CMU-MOSEI [48] A,V,T Dimensional 1 1
CH-SIMS [12] A,V,T Dimensional 1 1
CH-SIMS v2 [13] A,V,T Dimensional 1 1
VAM [49] A,V,T Dimensional 3 1
SEMAINE [14] A,V,T Dimensional 5 1
AFEW-VA [50] A,V,T Dimensional 2 1
SEWA[51] A,V,T Dimensional 3 1
MSP-Podcast [42] A Discrete 8 1
JL-Corpus [52] A Discrete 10 1
EmoDB [53] A Discrete 7 1
EMOVO [54] A Discrete 7 1
MESD [55] A Discrete 6 1
SFEW 2.0 [56] I Discrete 7 1
FER-2013 [57] I Discrete 7 1
EmotioNet [58] I Discrete 23 1
AffectNet [59] I Discrete 7 1
ExpW [60] I Discrete 7 1
RAF-DB [7] I Discrete 19 1$\sim$2
CMU-MOSEI [48] A,V,T Discrete 6 1
eNTERFACE [61] A,V,T Discrete 6 1
SAVEE [62] A,V,T Discrete 7 1
AFEW 7.0 [63] A,V,T Discrete 7 1
MAFW [64] A,V,T Discrete 11 1
DFEW [65] A,V,T Discrete 7 1
CREMA-D [66] A,V,T Discrete 6 1
MSP-IMPROV [15] A,V,T Discrete 4 1
RAVDESS [67] A,V,T Discrete 8 1
IEMOCAP [16] A,V,T Discrete 10 1
MELD [17] A,V,T Discrete 7 1
MC-EIU [68] A,V,T Discrete 7 1
MER2023 [9] A,V,T Discrete 6 1
MER2024 [18] A,V,T Discrete 6 1
@@
Appendix I Duration Distribution of OV-MERD
@@
Appendix J Details in Dataset Construction
@@
To generate pre-annotated visual clues, we sample three frames from each video and input them into GPT-4V. In this section, we discuss the rationale behind the choice of the number of sampled frames. Specifically, we categorize visual clues into two types: (1) visual clues with relatively long durations; and (2) visual clues with fast movements, such as eye movements, head movements, and micro-expressions. For the first type, since the duration of most videos is between 1 and 4 seconds (see Appendix I) and the video content is usually continuous with only minor differences between adjacent frames (see Appendix E), uniformly sampling three frames are sufficient to capture this information; For the second type, we observe that current MLLMs (including the GPT-4V used in this paper) struggle to capture these fast movements. Increasing the number of sampled frames does not address this issue. Previous research has also shown that GPT-4V cannot recognize micro-expressions [69]. To capture these fast movements, we employ multiple professional annotators to manually add this information.
@@
Appendix K Annotation Details
This section presents our annotation guidelines and the layout of the annotation platform. Our annotation process relies on the Label Studio [70] toolkit. As shown in Figure 2, there are two parts that require manual checking: 1) the pre-annotated acoustic and visual clues; 2) the merged open-vocabulary labels. To reduce subjective bias, we hire eight annotators who are experts in affective computing and familiar with the definitions of emotions. Additionally, we conduct two rounds of checks with no overlap among annotators in each round. Specifically, in the first round, we randomly select four annotators to check the clues and labels; in the second round, we merge the clues and labels reviewed by the first four annotators and ask another four annotators to perform a second round of checks. Ultimately, we find that these checked clues and labels are well-aligned with the video content.
@@
To illustrate which labels are removed or kept, we provide two examples, each with labels from multiple annotators. To ensure annotation quality, we hire professional annotators who are experts in affective computing and familiar with the definition of emotions. Some of these annotators are members of our team who specialize in affective computing. In Figure 17, as the character doesn’t know which doctor to see, most annotators provide labels such as confused or puzzled. Based on his tone and expression, some annotators further provide labels like anxious and serious. In Figure 18, most annotators notice his disapprove based on the textual content. Combining other modalities, some annotators further note his blame and accuse of what others are planning to do. From these examples, we can observe that these annotators provided relatively reliable labels. However, some annotators may focus only on the most relevant labels and overlook some details. To ensure the comprehensiveness of the annotation results, we merge the labels checked by four annotators. For example, in Figure 17, the final merged labels are troubled, focused, puzzled, anxious, worried, confused, and serious. In the next round, we invite another four annotators for a second check. Through this process, we can ensure that each preserved label is confirmed by at least one annotator in each round, thereby ensuring the comprehensiveness and accuracy of annotation results.
@@
Appendix L CLUE-Multi Analysis
@@
Besides discrete labels, MER2023 also provides relatively reliable valence scores. Therefore, we attempt to verify whether CLUE-Multi can used for valence estimation. Through experimental analysis, we observe that the PCC score between predictions and annotations can reach 0.88, indicating that CLUE-Multi also contains clues for dimensional emotion recognition.
@@
Appendix M Details of Language Impact Experiments
@@
Appendix N Cost of GPT-based Metrics
@@
Appendix O Emotion Wheel
@@
Appendix P Visualization of CLUE-M/A/T/V
@@
Appendix Q Details of CLUE-MLLM
@@
SECap [71] https://github.com/thuhcsi/SECap
SALMONN [19] https://github.com/bytedance/SALMONN
Qwen-Audio [72] https://github.com/QwenLM/Qwen-Audio
Otter [73] https://github.com/Luodian/Otter
OneLLM [74] https://github.com/csuhan/OneLLM
PandaGPT [75] https://github.com/yxuansu/PandaGPT
VideoChat [76] https://github.com/OpenGVLab/Ask-Anything/tree/main/video_chat
VideoChat2 [77] https://github.com/OpenGVLab/Ask-Anything/tree/main/video_chat2
Video-LLaMA [78] https://github.com/DAMO-NLP-SG/Video-LLaMA
Video-LLaVA [79] https://github.com/PKU-YuanGroup/Video-LLaVA
Video-ChatGPT [80] https://github.com/mbzuai-oryx/Video-ChatGPT
LLaMA-VID [81] https://github.com/dvlab-research/LLaMA-VID
mPLUG-Owl [82] https://github.com/X-PLUG/mPLUG-Owl
Chat-UniVi [83] https://github.com/PKU-YuanGroup/Chat-UniVi
GPT-4V [84] https://openai.com/
@@
Appendix R Relationship between Description Length and Label Numbers
@@
Appendix S GPT-based vs. EW-based Grouping
We explore the relationship between GPT-based and EW-based grouping strategies. We calculate the scores for different types of EW-based grouping strategies, with the average score denoted as M-avg. Table 12 reports $\mbox{F}_{\mbox{s}}$, as this metric is used for the final ranking, and we also compute the PCC scores between different metrics. We observe that the PCC score between M-avg and GPT is relatively high, indicating that EW-based metrics can serve as alternatives to GPT-based metrics.
Model GPT M1 M2 M3-W1 M3-W2
L1 L2 L1 L2
PCC score 0.887 0.857 0.911 0.940 0.913 0.942
Model M3-W3 M3-W4 M3-W5 M-avg
L1 L2 L1 L2 L1 L2
PCC score 0.904 0.927 0.899 0.922 0.885 0.894 0.942
Appendix T GPT-based vs. Matching-based Metrics

### v3

@@ Appendix D Related Work
MER has rapidly developed in recent years (Wu et al., 2014). Current research mainly focuses on building more efficient architectures to achieve higher accuracy on benchmark datasets (Sun et al., 2024). For example, Zadeh et al. (2017) proposed a tensor fusion network that addressed the MER task by leveraging interactions among unimodal, bimodal, and trimodal inputs. Tsai et al. (2019a) introduced a Transformer-based model that learned implicit alignment between different modalities and achieved promising results. Lian et al. (2024b) further established MERBench, involving various features, fusion strategies, and datasets. In emotion recognition, benchmark datasets usually limit the label space to basic emotions and use majority voting to determine the most likely one or more labels (Lian et al., 2023; Li et al., 2017). However, emotional categories extend far beyond basic emotions. Restricting the label space will inevitably overlook some nuanced emotions. To address this issue, we extend traditional MER to OV-MER, which allows for the prediction of any number and category of emotions.
@@
Its main goal is to identify categories beyond the annotated label space (Wu et al., 2024), which has been applied in various fields, such as object detection (Zareian et al., 2021), segmentation (Ghiasi et al., 2022), and scene understanding (Li et al., 2021). For example, the object detection dataset COCO (Lin et al., 2014) contains 80 categories, while objects in the real world are nearly infinite, highlighting the importance of open vocabulary learning. This paper makes the first attempt to address MER in an open-vocabulary manner. Compared to other tasks, MER is more difficult as it requires considering multimodal clues and subtle temporal variations.
Appendix E MER2023 Details
Our dataset utilizes videos sourced from MER2023 (Lian et al., 2023), and the use of this dataset has been consented to by the dataset owners. During the construction of this dataset, MER2023 employs a voice activity detection (VAD) tool to split videos based on the presence or absence of human speech. Subsequently, it uses a tool to measure speaker similarity and merge consecutive clips from the same speaker, thereby ensuring relatively complete content for each video clip. Afterward, multiple filters are applied to remove videos with inappropriate lengths or those with multiple speakers. Finally, most samples in MER2023 are single-person videos with relatively complete speech content.
Appendix F More Examples
Figures 8$\sim$10 provide more examples to visualize the difference between one-hot and OV labels. This paper uses emotion-related descriptions as a bridge for OV label extraction. Since the original video contains real people, we use DemoAI to remove personal information to address copyright concerns. Our OV-MERD dataset is derived from the MER2023 dataset (Lian et al., 2023) with further annotations. Therefore, for the original data, please download the MER2023 dataset.
@@ Appendix F More Examples
Appendix G Summary of Abbreviations
@@
Appendix H Dataset Comparison
@@
MSP-Podcast (Lotfian & Busso, 2017) A Dimensional 3 1
SST (Socher et al., 2013) T Dimensional 1 1
Cornell (Pang et al., 2002) T Dimensional 1 1
Large Movie (Maas et al., 2011) T Dimensional 1 1
ICT-MMMO (Wöllmer et al., 2013) A,V,T Dimensional 1 1
YouTube (Morency et al., 2011) A,V,T Dimensional 1 1
MOUD (Pérez-Rosas et al., 2013) A,V,T Dimensional 1 1
CMU-MOSI (Zadeh et al., 2017) A,V,T Dimensional 1 1
CMU-MOSEI (Zadeh et al., 2018b) A,V,T Dimensional 1 1
CH-SIMS (Yu et al., 2020) A,V,T Dimensional 1 1
CH-SIMS v2 (Liu et al., 2022b) A,V,T Dimensional 1 1
VAM (Grimm et al., 2008) A,V,T Dimensional 3 1
SEMAINE (McKeown et al., 2011) A,V,T Dimensional 5 1
AFEW-VA (Kossaifi et al., 2017) A,V,T Dimensional 2 1
SEWA(Kossaifi et al., 2019) A,V,T Dimensional 3 1
MSP-Podcast (Lotfian & Busso, 2017) A Discrete 8 1
JL-Corpus (James et al., 2018) A Discrete 10 1
EmoDB (Burkhardt et al., 2005) A Discrete 7 1
EMOVO (Costantini et al., 2014) A Discrete 7 1
MESD (Duville et al., 2021) A Discrete 6 1
SFEW 2.0 (Dhall et al., 2015) I Discrete 7 1
FER-2013 (Goodfellow et al., 2013) I Discrete 7 1
EmotioNet (Fabian Benitez-Quiroz et al., 2016) I Discrete 23 1
AffectNet (Mollahosseini et al., 2017) I Discrete 7 1
ExpW (Zhang et al., 2018) I Discrete 7 1
RAF-DB (Li et al., 2017) I Discrete 19 1$\sim$2
CMU-MOSEI (Zadeh et al., 2018b) A,V,T Discrete 6 1
eNTERFACE (Martin et al., 2006) A,V,T Discrete 6 1
SAVEE (Jackson & Haq, 2014) A,V,T Discrete 7 1
AFEW 7.0 (Dhall et al., 2017) A,V,T Discrete 7 1
MAFW (Liu et al., 2022a) A,V,T Discrete 11 1
DFEW (Jiang et al., 2020) A,V,T Discrete 7 1
CREMA-D (Cao et al., 2014) A,V,T Discrete 6 1
MSP-IMPROV (Busso et al., 2016) A,V,T Discrete 4 1
RAVDESS Livingstone & Russo (2018) A,V,T Discrete 8 1
IEMOCAP (Busso et al., 2008) A,V,T Discrete 10 1
MELD (Poria et al., 2019) A,V,T Discrete 7 1
MC-EIU (Liu et al., 2024) A,V,T Discrete 7 1
MER2023 (Lian et al., 2023) A,V,T Discrete 6 1
MER2024 (Lian et al., 2024a) A,V,T Discrete 6 1
@@
Appendix J Duration Distribution of OV-MERD
@@
Appendix K Details in Dataset Construction
@@
To generate pre-annotated visual clues, we sample three frames from each video and input them into GPT-4V. In this section, we discuss the rationale behind the choice of the number of sampled frames. Specifically, we categorize visual clues into two types: (1) visual clues with relatively long durations; and (2) visual clues with fast movements, such as eye movements, head movements, and micro-expressions. For the first type, since the duration of most videos is between 1 and 4 seconds (see Appendix J) and the video content is usually continuous with only minor differences between adjacent frames (see Appendix F), uniformly sampling three frames are sufficient to capture this information; For the second type, we observe that current MLLMs (including the GPT-4V used in this paper) struggle to capture these fast movements. Increasing the number of sampled frames does not address this issue. Previous research has also shown that GPT-4V cannot recognize micro-expressions (Lian et al., 2024c). To capture these fast movements, we employ multiple professional annotators to manually add this information.
@@
Appendix L Annotation Details
This section presents our annotation guidelines and the layout of the annotation platform. Our annotation process relies on the Label Studio (Tkachenko et al., 2020) toolkit. As shown in Figure 2, there are two parts that require manual checking: 1) the pre-annotated acoustic and visual clues; 2) the merged open-vocabulary labels. To reduce subjective bias, we hire eight annotators who are experts in affective computing and familiar with the definitions of emotions. To maintain high-quality annotations, all annotators must pass a rigorous preliminary test. This test evaluates their performance on 12 samples, each of which was previously annotated by five annotators with full agreement. Annotators who perform poorly are removed from the annotator pool. Additionally, we conduct two rounds of checks with no overlap among annotators in each round. Specifically, in the first round, we randomly select four annotators to check the clues and labels; in the second round, we merge the clues and labels reviewed by the first four annotators and ask another four annotators to perform a second round of checks. Ultimately, we find that these checked clues and labels are well-aligned with the video content.
@@
To illustrate which labels are removed or kept, we provide two examples, each with labels from multiple annotators. To ensure annotation quality, we hire professional annotators who are experts in affective computing and familiar with the definition of emotions. Some of these annotators are members of our team who specialize in affective computing. In Figure 17, as the character doesn’t know which doctor to see, most annotators provide labels such as confused or puzzled. Based on his tone and expression, some annotators further provide labels like anxious and serious. In Figure 18, most annotators notice his disapproval based on the textual content. Combining other modalities, some annotators further note his blame and accuse of what others are planning to do. From these examples, we can observe that these annotators provided relatively reliable labels. However, some annotators may focus only on the most relevant labels and overlook some details. To ensure the comprehensiveness of the annotation results, we merge the labels checked by four annotators. For example, in Figure 17, the final merged labels are troubled, focused, puzzled, anxious, worried, confused, and serious. In the next round, we invite another four annotators for a second check. Through this process, we can ensure that each preserved label is confirmed by at least one annotator in each round, thereby ensuring the comprehensiveness and accuracy of annotation results.
@@
Inter-annotator Agreement.
In this section, we calculate the inter-annotator agreement for two-round checks. Unlike the traditional single-label-based annotation method with a fixed label space, OV-MER employs a multi-label-based annotation method without a fixed label space. Therefore, we cannot directly compute the Kappa value between different annotators. To this end, we draw inspiration from Section N and utilize the Jaccard similarity coefficient to measure the inter-annotator agreement. Specifically, assume there are $N$ samples and $K$ annotators. For each pair of annotators $A_{m}$ and $A_{n}$, their annotation results for each sample $x_{i}$ are denoted as $Y_{m}^{i}$ and $Y_{n}^{i}$, respectively. Here, $Y_{m}^{i}$ and $Y_{n}^{i}$ contain a set of emotion labels. We calculate the agreement score between annotators $A_{m}$ and $A_{n}$ as:
$\mbox{Similarity}_{m,n}=\frac{1}{N}\sum_{i=1}^{N}\frac{|Y_{m}^{i}\cap Y_{n}^{i}|}{|Y_{m}^{i}\cup Y_{n}^{i}|}.$ (5)
In our annotation process, we hired 8 annotators and conducted two rounds of checks, with no overlap among annotators in each round. Table 8 presents the inter-annotator agreement for the first round, and Table 9 presents the inter-annotator agreement for the second round. We observe that through multi-round checks, the inter-annotator agreement gradually increases. These results demonstrate the necessity of multi-round checks, which help enhance label reliability.
Table 8: Inter-annotator agreement in the first round of annotation.
$A_{1}$ $A_{2}$ $A_{3}$ $A_{4}$
$A_{1}$ 1.00 0.57 0.47 0.51
$A_{2}$ 0.57 1.00 0.49 0.48
$A_{3}$ 0.47 0.49 1.00 0.46
$A_{4}$ 0.51 0.48 0.46 1.00
Table 9: Inter-annotator agreement in the second round of annotation.
$A_{5}$ $A_{6}$ $A_{7}$ $A_{8}$
$A_{5}$ 1.00 0.66 0.71 0.77
$A_{6}$ 0.66 1.00 0.64 0.67
$A_{7}$ 0.71 0.64 1.00 0.69
$A_{8}$ 0.77 0.67 0.69 1.00
Appendix M CLUE-Multi Analysis
@@
Besides discrete labels, MER2023 also provides relatively reliable valence scores. Therefore, we attempt to verify whether CLUE-Multi can be used for valence estimation. Through experimental analysis, we observe that the PCC score between predictions and annotations can reach 0.88, indicating that CLUE-Multi also contains clues for dimensional emotion recognition.
@@
Appendix N Details of Language Impact Experiments
@@
Appendix O Cost of GPT-based Metrics
@@
Appendix P Emotion Wheel
@@
Appendix Q Visualization of CLUE-M/A/T/V
@@
Appendix R Details of CLUE-MLLM
@@
SECap (Xu et al., 2024) https://github.com/thuhcsi/SECap
SALMONN (Tang et al., 2023) https://github.com/bytedance/SALMONN
Qwen-Audio (Chu et al., 2023) https://github.com/QwenLM/Qwen-Audio
Otter (Li et al., 2023a) https://github.com/Luodian/Otter
OneLLM (Han et al., 2024) https://github.com/csuhan/OneLLM
PandaGPT (Su et al., 2023) https://github.com/yxuansu/PandaGPT
VideoChat (Li et al., 2023b) https://github.com/OpenGVLab/Ask-Anything/tree/main/video_chat
VideoChat2 (Li et al., 2024a) https://github.com/OpenGVLab/Ask-Anything/tree/main/video_chat2
Video-LLaMA (Zhang et al., 2023) https://github.com/DAMO-NLP-SG/Video-LLaMA
Video-LLaVA (Lin et al., 2024) https://github.com/PKU-YuanGroup/Video-LLaVA
Video-ChatGPT (Maaz et al., 2024) https://github.com/mbzuai-oryx/Video-ChatGPT
LLaMA-VID (Li et al., 2024b) https://github.com/dvlab-research/LLaMA-VID
mPLUG-Owl (Ye et al., 2023) https://github.com/X-PLUG/mPLUG-Owl
Chat-UniVi (Jin et al., 2024) https://github.com/PKU-YuanGroup/Chat-UniVi
GPT-4V (OpenAI, 2023) https://openai.com/
@@
Appendix S Cross-linguistic Correlation
In Table 14, we leverage the results from Table 2 to compute PCC scores between the English and Chinese results for each metric. Experimental results demonstrate that all metrics exhibit strong cross-linguistic correlations.
Table 14: PCC between the English and Chinese results for each metric.
$\mbox{F}_{\mbox{s}}$ $\mbox{Precision}_{\mbox{s}}$ $\mbox{Recall}_{\mbox{s}}$
PCC scores 0.9896 0.9738 0.9817
Appendix T Relationship between Description Length and Label Numbers
@@
Appendix U Performance of Discriminative MER Methods
This paper primarily focuses on MLLM-based generative models for OV-MER. Traditional discriminative methods are not applied due to fundamental differences in our experimental setup. Specifically, discriminative methods require identical label spaces $\mathcal{Y}$ for both training and testing sets. They cannot predict unseen emotions, i.e., $y\notin\mathcal{Y}$. However, we use an open-vocabulary annotation manner in OV-MER, which inherently cannot guarantee alignment between training and testing label spaces (i.e., $\mathcal{Y}_{train}\neq\mathcal{Y}_{test}$). MLLM-based generative methods offer greater flexibility in emotion prediction, making them better suited for our task. Consequently, we primarily leverage MLLM-based solutions. If forced to use discriminative approaches, these models could only predict labels within their training label space $\mathcal{Y}_{train}$.
In Table 15, we follow the zero-shot experimental setup commonly used in generative models and report results for discriminative models. Specifically, we train on the IEMOCAP (Busso et al., 2008) (or MELD (Poria et al., 2019)) dataset and evaluate on OV-MERD. For discriminative models, we use CLIP-Large for visual features, HUBERT-Large for acoustic features, and Baichuan-13B for lexical features, comparing the performance of different classifiers. In this table, the “Attention” model refers to a foundation model architecture in MERBench (Lian et al., 2024b). Specifically, let $f_{i}^{a}\in\mathbb{R}^{d_{a}}$, $f_{i}^{v}\in\mathbb{R}^{d_{v}}$, and $f_{i}^{l}\in\mathbb{R}^{d_{l}}$ denote the acoustic, visual, and lexical features for a sample $x_{i}$, respectively. This model first converts all inputs into the same dimension and then computes importance scores $\alpha_{i}$ for each modality. Subsequently, it employs weighted fusion to obtain multimodal features $z_{i}$, which are utilized for emotion prediction.
$h_{i}^{m}=\mbox{ReLU}\left(f_{i}^{m}W_{m}^{h}+b_{m}^{h}\right),m\in\{a,l,v\},$ (8)
$h_{i}=\mbox{Concat}\left(h_{i}^{a},h_{i}^{l},h_{i}^{v}\right),$ (9)
$\alpha_{i}=\mbox{softmax}\left(h_{i}^{T}W_{\alpha}+b_{\alpha}\right),$ (10)
$z_{i}=h_{i}\alpha_{i}.$ (11)
Here, $W_{m}^{h}\in\mathbb{R}^{d_{m}\times h}$, $b_{m}^{h}\in\mathbb{R}^{h}$, $W_{\alpha}\in\mathbb{R}^{h\times 1}$, and $b_{\alpha}\in\mathbb{R}^{3}$ are trainable parameters. For the output, we have $h_{i}^{m}\in\mathbb{R}^{h}$, $h_{i}\in\mathbb{R}^{h\times 3}$, $\alpha_{i}\in\mathbb{R}^{3\times 1}$, and $z_{i}\in\mathbb{R}^{h}$. Experimental results in Table 15 show that while discriminative models can be adapted to solve OV-MER, they generally perform worse than MLLM-based generative models.
Table 15: Zero-shot performance of traditional discriminative models and MLLM-based generative models.
Model M3-W1 M3-W2 M3-W3 M3-W4 M3-W5
L1 L2 L1 L2 L1 L2 L1 L2 L1 L2
Traditional Discriminative Models
MELD + MFM (Tsai et al., 2019b) 22.28 13.51 21.77 13.51 19.59 17.67 22.10 18.20 16.72 14.82
MELD + MISA (Hazarika et al., 2020) 28.72 21.75 27.59 22.43 34.31 28.50 26.19 21.80 34.79 29.24
MELD + GMFN (Zadeh et al., 2018b) 34.28 22.16 33.77 22.47 32.40 29.16 33.43 28.18 29.43 26.50
MELD + MFN (Zadeh et al., 2018a) 31.19 21.57 30.66 21.66 31.26 28.02 32.42 25.54 27.97 24.89
MELD + MulT (Tsai et al., 2019a) 30.74 17.76 30.67 18.45 28.08 23.58 29.89 23.68 24.72 20.79
MELD + LMF (Liu et al., 2018) 41.47 27.70 40.86 28.43 42.29 37.36 38.54 32.83 40.05 35.16
MELD + TFN (Zadeh et al., 2017) 31.91 20.54 31.41 20.56 31.15 26.75 28.41 23.81 29.68 25.36
MELD + Attention (Lian et al., 2024b) 33.61 23.16 32.27 23.42 35.17 30.41 30.88 25.75 33.72 29.53
IEMOCAP + MFM (Tsai et al., 2019b) 45.46 32.86 47.55 33.12 46.37 39.90 43.03 36.97 43.97 39.28
IEMOCAP + MISA (Hazarika et al., 2020) 49.14 35.98 48.80 36.53 48.66 43.86 47.31 39.82 48.21 43.37
IEMOCAP + GMFN (Zadeh et al., 2018b) 49.35 35.85 49.57 36.09 49.18 43.29 46.72 39.28 47.30 42.71
IEMOCAP + MFN (Zadeh et al., 2018a) 50.56 36.82 50.86 36.72 49.97 44.70 48.69 40.55 48.97 44.11
IEMOCAP + MulT (Tsai et al., 2019a) 42.67 30.27 43.50 30.79 42.10 37.21 40.75 34.31 41.00 36.55
IEMOCAP + LMF (Liu et al., 2018) 46.34 32.44 46.42 32.94 44.19 39.22 44.23 36.78 43.57 38.57
IEMOCAP + TFN (Zadeh et al., 2017) 46.13 33.45 46.66 33.91 46.27 41.27 42.31 35.95 45.82 40.69
IEMOCAP + Attention (Lian et al., 2024b) 45.64 32.23 46.18 32.31 44.42 39.23 43.40 36.67 43.65 38.49
MLLM-based Generative Models
Qwen-Audio (Chu et al., 2023) 43.85 26.60 41.52 26.68 39.46 30.65 36.64 27.33 35.89 29.66
Otter (Li et al., 2023a) 49.75 33.50 49.93 33.04 51.03 37.12 47.54 34.77 50.51 35.54
Video-LLaMA (Zhang et al., 2023) 52.90 36.08 53.60 35.33 47.50 36.50 52.97 35.78 46.39 34.77
VideoChat (Li et al., 2023b) 47.79 32.64 47.76 32.14 46.78 34.37 49.53 32.82 45.93 32.85
SECap (Xu et al., 2024) 52.26 37.55 52.11 37.71 50.77 40.49 50.43 38.21 49.97 40.25
Video-LLaVA (Lin et al., 2024) 54.65 37.65 54.54 38.25 52.29 40.58 52.45 39.91 52.97 39.69
SALMONN (Tang et al., 2023) 54.90 38.93 54.29 37.79 56.25 43.01 50.53 38.54 53.65 42.09
VideoChat2 (Li et al., 2024a) 52.38 36.44 53.56 36.91 52.14 40.57 50.63 39.64 51.37 39.89
Video-ChatGPT (Maaz et al., 2024) 57.66 41.48 57.37 40.95 55.50 44.15 55.24 42.42 52.93 41.54
LLaMA-VID (Li et al., 2024b) 56.59 41.22 57.49 40.39 55.12 44.06 56.62 42.42 53.03 41.65
mPLUG-Owl (Ye et al., 2023) 57.60 41.32 56.32 40.83 55.67 43.71 55.06 40.67 54.44 42.00
Chat-UniVi (Jin et al., 2024) 57.00 42.25 57.50 42.43 56.80 45.66 55.86 41.97 55.81 43.61
Appendix V GPT-based vs. Matching-based Metrics